Suppose that the coefficient of the quadratic term and the coefficient of the linear term are equal. If the quadratic has a double root , what is the ratio of the coefficient the quadratic term to the constant term
Found 2 solutions by rothauserc, ikleyn:
Answer by rothauserc(4718)

The general form for a quadratic is
:
ax^2 +bx +c
:
It is given that a = c
:
Note the discriminant b^2 -4ac = 0 when a quadratic has a double root
:
a^2 -4ac = 0
:
Since a and c are not = 0(why?), then
:
a -4c = 0
:
a = 4c
:
a/c = 4
:

Answer by ikleyn(52780)

.

In the post by @rothauserc, the third line should be read as


    a = b.


The rest of the post, the solution itself and the answer are correct.